OpenStudy (jakyfraze):

can someone please help me and show me how to do this please ;"( Simplify and solve this equation: 4m + 9 + 5m – 12 = 42. A. m = 5 B. m = 4 C. m = –5 D. m = -4

OpenStudy (anonymous):

So the way it works is you have to plug it in. so start with 4. 16+9+20-12 that is 45-12 which is 33. So no. 5 would be 20 + 9 + 25 - 12 which is 42! So the answer is 5. Just remember, plug in the numbers, and everything would be great! Lmk if you need anything else

OpenStudy (jakyfraze):

were did you get this from 4. 16+9+20-12?

OpenStudy (jakyfraze):

what happened to this 4m + 9 + 5m – 12 = 42.

OpenStudy (anonymous):

so it is 4m so that is 4x4 which is 16. And 5m would be 5 x 4 which is 20

OpenStudy (anonymous):

I just plugged in the #s. Sorry!

OpenStudy (anonymous):

4m + 9 + 5m – 12 = 42. that is the base. so 4m = 4 x the #, so you replace 4m with the answer

OpenStudy (jakyfraze):

I'm really sorry. i don't understand :(.

OpenStudy (anonymous):

Ok, so 4m is 4 times of the variable. So m is the variable. So what you are doing is that you replace the variable with the number, so 4m would be 4 x 4 if you are replacing m with 4. So m would be 5 if you use 5 as the variable. And hey, it is fine!

OpenStudy (jakyfraze):

thank you :'). so how is 4m 4x4 but 5m is not 5x5

OpenStudy (anonymous):

I am using 4m as an example from the problem! so up there, there is 4m and 5m. So if m=5 then there would be 4 x 5 and 5 x 5 because it is 4m/5m. The 4/5m means that you are multiplying the number times the variable

OpenStudy (anonymous):

Do you understand or no? Please let me know!

OpenStudy (jakyfraze):

so what do you mean by 4 x 5 and 5 x 5?

OpenStudy (jakyfraze):

im jus trying to understand

OpenStudy (anonymous):

so it is 4m which is 4 times the variable. the letter is the variable. So 4x5 is 4m and m=4 so when a variable is connected to the #, that means to multiply it. So when you do 4m it = 4 x4. and so up there, 5m would be 5 times the variable. And since m is the variable, m would = one of the 4 answer choices. Just plug them in, and it would be fine. so 4m = 4 x the variable

OpenStudy (anonymous):

LMK if you need more help!

OpenStudy (jakyfraze):

I'm really sorry but i really don't understand at all

OpenStudy (anonymous):

Really it is no problem. So tell me, what is it that you don't understand? Is it how the 4m = 4 x the #, or is it the m = a # or? I need to know so I can help!

OpenStudy (jakyfraze):

so it is 4m which is 4 times the variable. the letter is the variable. So 4x5 is 4m and m=4 so when a variable is connected to the #, that means to multiply it. So when you do 4m it = 4 x4. and so up there, 5m would be 5 times the variable. And since m is the variable, m would = one of the 4 answer choices. Just plug them in, and it would be fine. so 4m = 4 x the variable

OpenStudy (jakyfraze):

what i don't understand is the whole entire question and the m thing you talking about. please don't think I'm trolling you or being funny.

OpenStudy (anonymous):

@jakyfraze what you said above was absolutely correct! so normally x is the normal variable, but in this case, they are using m. so you do m x the # and then add and subtract the numbers above in the correct order (PEMDAS) and then you get an answer, and if it matches, it is right!

OpenStudy (anonymous):

And I don't think you are trolling don't worry!

OpenStudy (jakyfraze):

oh :') thank you. can you please do a step by step on how you got the answer please. i really want to learn this. thank you so much trying to explain it to me :)

OpenStudy (anonymous):

Ok! So step one find the variables and such. So that in this case is m. And up there, the answer choices are 4, 5, -4, -5. so what you have to do is replace m with one of the answer choices. So in this case, it is 4m and 5m, so you replace them with 4 x 4,5,-4, or -5. Only one though. So let's use 4. 4m + 9 + 5m – 12 = 42. So replace 4m with 4 x 4 and 5m with 5 x 5. So once you have that, you just solve it. so 16+9+20-12. and that is 33. so let's redo it with 5. 20+9+25-12, and that is 42!

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

I guess not! I hope you understand! Have a great day and good luck!

OpenStudy (jakyfraze):

@WoahItsJian sorry i feel asleep :(

OpenStudy (jakyfraze):

@WoahItsJian 5 x 5= 25. you said 20.

OpenStudy (jakyfraze):

@Mehek14 are you good at math?

Mehek (mehek14):

add like terms

OpenStudy (jakyfraze):

can you explain to me how you would solve this or how to solve it

Mehek (mehek14):

4m + 5m = 9m 9 - 12 = -3 9m - 3 = 42 add 3 9m = 45 divide by 9 m = 5

OpenStudy (jakyfraze):

@Mehek14 why add three? are suppose to only add the middle to the right side?

Mehek (mehek14):

you want to get m by itself so you move everything to the other side when you move something, you do the opposite operation -3 becomes +3

OpenStudy (jakyfraze):

so you only do this when the number is -3?

OpenStudy (jakyfraze):

@Mehek14 right?

Mehek (mehek14):

no to anything moving to the other side

OpenStudy (jakyfraze):

what moves to the right side?

Mehek (mehek14):

like how you had 9m = 45 to get 9 to the other side, you had to do the opposite 9 was being multiplied to m to move it to the other side, i divided 9

OpenStudy (jakyfraze):

@Mehek14 can you give me a problem to solve please? i want to see if i got it

Mehek (mehek14):

ok